OpenStudy (anonymous):

I'm so behind in Algebra right now. Could anyone help me with these three questions? 1. What is the slope of a line that passes through (-2, -3) and (1, 1) 2. What is an equation for the line with slope 2/3 and y-intercept 9 3. What is an equation in slope-intercept form for the line that passes through the points (1, -3) and (3, 1)

OpenStudy (hyper):

1) Slope of a line is given by : m = \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] Therefore, the slope is : m = (1--3)/(1--2) = 4/3

OpenStudy (hyper):

2) Equation of a line in slope-intercept form is : y = mx + b, where m = slope and b = y-intercept. Try and sub in the number to get the line ...